Suppose that you have a reflection diffraction grating with n= 140 lines per millimeter. Light from a sodium lamp passes through the grating and is diffracted onto a distant screen.

A. Two visible lines in the sodium spectrum have wavelengths 498 nm and 569 nm. What is the angular separation Δθ of the first maxima of these spectral lines generated by this diffraction grating?

B. How wide does this grating need to be to allow you to resolve the two lines 589.00 and 589.59 nanometers, which are a well known pair of lines for sodium, in the second order (m=2)?
Answer #1

A)

d*sin(theta1) = m

(1/140000)*sin(theta1) = 1*498*10^-9

theta1 = 3.9979 deg.

d*sin(theta2) = m

(1/140000)*sin(theta2) = 1*569*10^-9

theta2 = 4.569 deg.

(theta) theta2 - theta1

theta = 4.569 - 3.9979

theta = 0.5711 deg

b)

Chromatic Resolving power,

R = /

R = 589 / (589.59 - 589)

R = 589 / 0.59

/ = m*(width)*n

589 / 0.59 = 2*(width)*140000

width = 3.565 mm